Find the complete integral of partial differential equations z+xp-x2yq2-x3pq =0


1
Expert's answer
2021-01-20T13:32:04-0500

f(x,y,z,p,q)=xpq+yq2-1=0"\frac{dp}{\frac{df}{dx}+p\frac{df}{dz}}=\frac{dq}{\frac{df}{dy}+q\frac{df}{dz}}=\frac{dz}{-p\frac{df}{dp}-q\frac{df}{dq}}=\frac{dx}{-\frac{df}{dp}}=\frac{dy}{-\frac{df}{dq}}"

"\frac{dp}{p-2xyq^2-3x^2pq+p}=\frac{dq}{-x^2q^2+q}=\frac{dz}{-p(x-x^3q)-q(2qyx^2-x^3p)}=\frac{dx}{-x+x^3q}=\frac{dy}{2qyx^2+x^3p}"


"\frac{\frac{dq}{q}}{1-qx^2}=\frac{\frac{dx}{x}}{x^2q-1}"

"\frac{dq}{q}+ \frac{dx}{x}=0"

ln(q)+ln(x)=a

qx=ea

q=ea/x


"\frac{ \frac{dp}{qp}+ \frac{dy}{xp}}{\frac{2}{q} - \frac{2xyq}{p}-3x^2+ \frac{2qyx}{p}+ x^2}=\frac{ \frac{dp}{qp}+ \frac{dy}{xp}}{\frac{2}{q} -2x^2}=0"

"\frac{dp}{q}+ \frac{dy}{x}=0"

"\frac{p}{q}=- \frac{y}{x} + b"

"\frac{e^ap}{x}=- \frac{y}{x} + b"

p=-yb/ea

dz=pdx+qdy=-yb/eadx+ea/xdy

z=-ybxea+eay/x+c=eayx (1/x2 - b)+c